mikewwe13:

A right triangle has side lengths that measure 4 units, 5 units, and √41 units. Is the value of the perimeter of the triangle a rational or irrational number? Is the value of the area of the triangle a rational or irrational number? Select from the drop-down menus to correctly complete the sentences. The value of the perimeter of the triangle is ____________ number. a rational or an irrational The value of the area of the triangle is ____________ number. a rational or irrational. Part B A square has side lengths that measure √5 units. Is the value of the perimeter of the square a rational or irrational number? Is the value of the area of the square a rational or irrational number? Select from the drop-down menus to correctly complete the sentences. The value of the perimeter of the square is ____________ number. a rational or an irrational The value of the area of the square is ___________ number. a rational or an irrational

Ultrilliam:

@Arcadiouse Can you help with this?

Arcadiouse:

I am currently helping one in Science but I can attempt in a bit I guess.

Arcadiouse:

I don't mind.

mikewwe13:

you know this ?

Arcadiouse:

Yeah, I have somewhat of an understanding, I am working on science problems and your problem at the same time, so give me a minute and I can help.

Arcadiouse:

Part A: Rational, Rational. That is what I got.

Arcadiouse:

Part B both are also rational. I swear this was easy but its odd that all four answers are rational.

mikewwe13:

are you sure ?

Arcadiouse:

I did my math multiple times, hold on ill do it again.

Arcadiouse:

Haha, Im wrong.

Arcadiouse:

Sorry dude, I would wait for @Vocaloid he does this type better than me. I have't done this math in forever.

mikewwe13:

do you know where vocaloid at ?

Ultrilliam:

For the love of... You guys know vocaloid is female right?

mikewwe13:

no

mikewwe13:

The area, A, of a trapezoid-shaped garden is given by the equation A = 1/2w(b + c) Hector knows the area of the garden, the width (w), and the length of one of the parallel sides (b). He wants to find the length of the other parallel side, c. Which of these shows the equation to find the length of the other parallel side, c? A. c = A/2w - b B. c = 2A/w - b C. c = 2A/bw D. c = A/2bw

Ultrilliam:

Well now you do.

Arcadiouse:

I didn't.

mikewwe13:

I DIDN'T KNOW

Ultrilliam:

Oh good Sam is viewing the question, @Sam any chance you can help them? x'D

Arcadiouse:

Yeah, I kinda forgot some of this. So I tried my best then realized I had my irrationals and rationals mixed up :(.

Arcadiouse:

And this is why I tend to stay out of math unless its math i ace lol.

Sam:

Oh wait you don't have to solve a trapezoid, you just have to mess around with the equation given

Sam:

You've got \[A = \frac{1}{2}w(b + c)\] So arranging this equation First multiply both sides by 2, \[2A=w(b+c)\] Divide both sides by w, \[\frac{2A}{w}=b+c\] Minus both sides by b and there's your c isolated

mikewwe13:

so it's B ?

mikewwe13:

@Sam